OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

gvgvh

OpenStudy (abb0t):

First, find the slope: \(\sf \color{blue}{slope} = m =\color{red}{\frac{y_2-y_1}{x_2-x_1}}\) it is also the change in points Once you have that, plug it into your point-slope formula: \(\sf \color{orange}{y-y_1=m(x-x_1)}\) where you can pick \(any\) of the two points you have. they will give you the SAME answer.

OpenStudy (abb0t):

You can leave it as is, or change it to a line, which is in the form: \(\sf \color{green}{y=mx+b}\)

OpenStudy (abb0t):

Can you finish it now?

OpenStudy (anonymous):

I got y+1=mx-7. Is that right?

OpenStudy (anonymous):

My assignment says the equation starts y=

OpenStudy (abb0t):

No. It's incorrect. Your first step is to find the slope. What is your slope?

OpenStudy (anonymous):

\[\frac{ -1 }{ 7}\]

OpenStudy (abb0t):

I see what you did wrong, but you picked the wrong points, I think you used your slope as your points. But your slope is just your slope. \(\sf \color{red}{m}\) is your slope. And it can be a fraction. It simply means rise over run: \(\sf \color{orange}{\frac{rise}{run}}\)

OpenStudy (abb0t):

Yes, correct. That's your slope. Now, plug it into: \(\sf \color{orange}{y-y_1=m(x-x_1)}\)

OpenStudy (abb0t):

for y\(_1\) and \(\sf x_1\) pick \(\sf \color{red}{one}\) of the two points you have. It can be either E or F

OpenStudy (anonymous):

\[y=\frac{ -1 }{ 7 }x\] is this right?

OpenStudy (abb0t):

So you have: \(\sf \color{green}{y-1=-\frac{1}{7}(x-5)}\) Can you solve it now, using algebra, to rearrange it into the form: \(\sf \color{red}{y=mx+b}\)

OpenStudy (anonymous):

\[y=\frac{ -1 }{ 7 }+4\] I think I got it this time